# YITP-24-61

### v2

@@ Abstract
Discussion. We studied a possibility of extending the theorem [3] to the case where the $O(4)$-symmetric instanton solution exhibits a singular behavior at $\rho=0$. In particular, we considered the case in which the solution does not satisfy Coleman’s condition (3). We further investigated a small anisotropic deformation around the singular $O(4)$-symmetric instanton. We proposed a concrete example with a piecewise potential given by (19). We then analyzed the solutions for $\bar{\Phi}$, that is singular at $\rho=0$ and goes to false vacuum as $\rho\rightarrow\infty$. Finally, we studied the dynamics of small anisotropic deformation and found that there exists a small and regular deformation around the singular $O(4)$-symmetric solution, which does not contribute to the total bounce action. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons, and gives rise to the possibility of the existence of a non-$O(4)$-symmetric instanton with a lower action. This opens several possibilities to study other cases, where the Coleman’s theorem may not apply. It would be interesting to explore whether the so-called pseudo bounces [10, 14] give lower or higher action than the one in our case, and study their implications.*10*10*10It could be that some off-shell configurations lead to lower actions, but they may correspond to subcritical bubbles. We leave this investigation to future work. Moreover, it is of interest to extend our analysis to the case with gravity. In connection with this, see e.g. [18, 19, 20, 21, 22, 23]. In our separate work [16] we will discuss other types of singular behaviors in addition to logarithmic divergence, and the generic form of the instanton potential.
Acknowledgements. We thank S. Mukohyama, N. Oshita, R. Saito and K. Takahashi for useful discussions. We thank J. R. Espinosa for pointing out an error in the first version of the paper. This work is supported in part by JSPS KAKENHI Nos. JP20H05853 and JP24K00624. M.S. and V.Y. are supported by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. Y.Z. is supported by the Fundamental Research Funds for the Central Universities, and by the Project 12475060 and 12047503 supported by NSFC, Project 24ZR1472400 sponsored by Natural Science Foundation of Shanghai, and also by China Scholarship Council.

### v3

@@ Abstract
Discussion. We studied a possibility of extending the theorem Coleman:1977th to the case where the $O(4)$-symmetric instanton solution exhibits a singular behavior at $\rho=0$. In particular, we considered the case in which the solution does not satisfy Coleman’s condition (3). We further investigated a small anisotropic deformation around the singular $O(4)$-symmetric instanton. We proposed a concrete example with a piecewise potential given by (19). In our separate work Sasaki:2025nzy we discuss the generic form of the instanton potential, and other types of singular behaviors in addition to logarithmic divergence. We then analyzed the solutions for $\bar{\Phi}$, that is singular at $\rho=0$ and goes to false vacuum as $\rho\rightarrow\infty$. Finally, we studied the dynamics of small anisotropic deformation and found that there exists a small and regular deformation around the singular $O(4)$-symmetric solution, which does not contribute to the total bounce action. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons, and gives rise to the possibility of the existence of a non-$O(4)$-symmetric instanton with a lower action. To clarify this point, it is necessary to make a non-linear analysis, which is a future issue. In any case, our result opens several possibilities to study other cases, where the Coleman’s theorem may not apply.
First of all, it could be that the quantum corrections to a singular instanton with finite action diverge. This issue also certainly deserves further study. Second, it would be interesting to explore whether the so-called pseudo bounces Espinosa:2019hbm; Espinosa:2021qeo give lower or higher action than the one in our case, and study their implications.*9*9*9It could be that some off-shell configurations lead to lower actions, but they may correspond to subcritical bubbles. Moreover, as discussed in a paragraph above Eq. (4), it is necessary to extend our analysis to the case with gravity. The results presented here may therefore be regarded as a first step toward a sophisticated study, in which gravity plays an important role. In connection with this, see e.g. Cohn:1998et; Gratton:1999ya; Dunne:2006bt; Ai:2023yce; Oshita:2023pwr; Oshita:2021aux.
Acknowledgements. We thank S. Mukohyama, N. Oshita, R. Saito and K. Takahashi for useful discussions. We thank J. R. Espinosa for pointing out an error in the first version of the paper. This work is supported in part by JSPS KAKENHI Nos. JP20H05853 and JP24K00624. M.S. and V.Y. are supported by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. V.Y. is supported in part by grants for development of new faculty staff, Ratchadaphiseksomphot Fund, Chulalongkorn University and by the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation Grant No. B39G680009. Y.Z. is supported by the Fundamental Research Funds for the Central Universities, and by the Project 12475060 and 12047503 supported by NSFC, Project 24ZR1472400 sponsored by Natural Science Foundation of Shanghai.